# Random Pareto front surfaces

### v1

@@ 4.4 Probability of domination
The probability that a vector or a set of vectors is Pareto optimal is given by the probability of domination. This probability can also be interpreted as a generalisation of the survival function for Pareto front surface distributions. By Lemma 3.1, we can write the probability of domination in term of the projected lengths
@@
for any set of vectors $Y\subseteq\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$, where $\land$ denotes the logical and operation. In practice, this probability is an intractable quantity. Nevertheless, if we can sample a collection of random Pareto front surfaces, then we can easily estimate it by using a simple Monte Carlo average: $\mathbb{P}[I(\boldsymbol{\omega})]\approx\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}[I(\boldsymbol{\omega}_{n})]$, where $I(\boldsymbol{\omega})\in\{0,1\}$ denotes a random boolean variable and $\boldsymbol{\omega}_{n}\in\Omega$ denotes some independent samples of the random parameter. In the left plot of Figure 8, we visualise the contours associated with the estimated marginal domination probabilities for the Pareto front surface distribution described in Figure 7.
@@ 4.5 Probability of deviation
In some cases, we might be interested in computing the probability that a vector or a set of vectors lies between two potentially random Pareto front surfaces. We refer to this quantity as the probability of deviation. By using Lemma 3.1, we can write the probability of deviation in term of the projected lengths
@@
for any random Pareto front surfaces $A^{*}(\boldsymbol{\omega}),B^{*}(\boldsymbol{\omega})\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$ and any set vectors $Y\subseteq\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$. As with the domination probabilities in (17), this quantity can be easily estimated using a Monte Carlo average. In Figure 8, we give an illustration of the estimated marginal deviation probabilities when $A^{*}(\boldsymbol{\omega})=Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega})$ is a random Pareto front surface and $B^{*}(\boldsymbol{\omega})=\mathbb{E}_{\boldsymbol{\omega}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega})]$ is a candidate for the expectation which is deterministic.
@@ 4.6 Vorob’ev statistics
Random set theory (Molchanov, 2005) gives us a way to extend standard concepts from probability and statistics to work over the space of closed sets. Some key ideas from this field have been used in earlier work in order to define some Pareto front surface statistics (Grunert da Fonseca and Fonseca, 2010; Binois et al., 2015a). Precisely, these ideas have focussed on defining probabilistic concepts over the space of domination regions $\mathcal{D}_{\preceq}:=\{\mathbb{D}_{\preceq}[A]\subset\mathbb{R}^{M}:A\subset\mathbb{R}^{M}\}$. In the following, we will review the core ideas within these works and show how they relate to our framework. To increase the generality of our discussion, we will consider working in the more general space of truncated domination regions $\mathcal{D}_{\preceq,\boldsymbol{\eta}}:=\{\mathbb{D}_{\preceq,\boldsymbol{\eta}}[A]\subset\mathbb{R}^{M}:A\subset\mathbb{R}^{M}\}$.
@@
In random set theory, one often relies on the concept of a coverage function, which gives us the probability that an element lies within a random closed set. When working over the space of truncated domination regions $\mathcal{D}_{\preceq,\boldsymbol{\eta}}$, this coverage function is simply given by the probability of domination (17): $\mathbb{P}[\mathbf{y}\in\mathbb{D}_{\preceq,\boldsymbol{\eta}}[Y^{*}_{\boldsymbol{\eta},f}]]$ for any vector $\mathbf{y}\in\mathbb{R}^{M}$. Equipped with this coverage function, we can now define many different types of probabilistic concepts by leveraging ideas from random set theory. Notably, the existing works referenced above have primarily focussed on extending the Vorob’ev concepts of random set theory (Molchanov, 2005, Section 2.2), which we describe below.
@@
In addition to proposing the Vorob’ev $\alpha$-quantile, Grunert da Fonseca and Fonseca (2010) proposed using the median quantile ($\alpha=0.5$) as a potential candidate for the mean Pareto front surface. In contrast Binois et al. (2015a) proposed using the Vorob’ev expectation as a candidate for the mean. The Vorob’ev expectation is defined as the $\alpha^{*}$-level Vorob’ev quantile whose hypervolume is the closest to the expected hypervolume of the random Pareto front surface. Precisely, $\mathbb{E}^{\text{Vorob'ev}}_{\boldsymbol{\omega}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega})]:=\mathcal{Q}_{\boldsymbol{\omega}}^{\text{Vorob'ev}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega}),\alpha^{*}]$, where $\alpha^{*}$ satisfies the inequality
$\nu[\mathcal{Q}^{\text{Vorob'ev}}_{\boldsymbol{\omega}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega}),\alpha]]\leq V_{\boldsymbol{\eta},f}\leq\nu[\mathcal{Q}^{\text{Vorob'ev}}_{\boldsymbol{\omega}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega}),\alpha^{*}]]$ (21)
@@
Analogous to how the traditional scalar-valued expectation minimises the variance, the Vorob’ev mean is known to minimise a quantity known as the Vorob’ev deviation. In the work by Binois et al. (2015a), they defined the Vorob’ev deviation of a Pareto front surface distribution as the expected hypervolume of the symmetric difference between the Pareto front surface and the considered set. Equivalently, the quantity can be written as the expected hypervolume distance (9) between these two sets, that is
@@
for any measurable set $A\in\mathcal{M}(\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}])$ lying in the truncated space. As shown in the work by Molchanov (2005, Section 2.2), the Vorob’ev mean is a minimiser to the Vorob’ev deviation over the space of measurable sets $A\in\mathcal{M}(\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}])$ subject to the condition that $\nu[A]=V_{\boldsymbol{\eta},f}$. That is,
@@
As a direct consequence of Proposition 4.2, the positive isolines of the Vorob’ev based Pareto front surfaces are all valid Pareto front surfaces because they are just instances of the quantile front (15). This result was taking for granted in the original works, but we have shown that it follows quite naturally from our polar parametrisation result.
@@
The Vorob’ev statistics above can be estimated using the empirical quantiles as we described in Section 4.3. For the Vorob’ev mean front, we have to additionally compute the parameter $\alpha^{*}$ which satisfies the hypervolume condition (21). To approximately solve this problem, Binois et al. (2015a) proposed using an iterative bisection scheme. Notably, this procedure can be very expensive because it requires evaluating the hypervolume of the quantiles several times. On the other hand, our definition of the expectation is much cheaper and simpler to estimate. In Figure 8, we present a visual comparison between the estimated Vorob’ev mean front and our sample mean front for a simple two-dimensional example. In this example, there does not appear to be any meaningful difference between these two Pareto front surfaces.
@@
4.7 Decision-theoretic construction
Many standard concepts in probability, concerning real-valued random variables, can be constructed under the decision-theoretic perspective in which we optimise a certain scoring rule (Gneiting, 2011). Remarkably, our work on the polar parametrisation of the Pareto front surface gives us an easy way to generalise this decision-theoretic construction to the space of Pareto front surface distributions. In particular, we can recover many of the Pareto front surface statistics introduced in the previous sections as being equal to a functional $F_{\boldsymbol{\omega}}$, which minimises some expected frontier loss (8),
@@ 5 Applications
Many of the concepts and results which we describe in this work are general and can be applied in any scenario where we want to infer or quantify properties about a random Pareto front surface. In this section, we identify a few concrete applications of these ideas which might be of interest for practitioners and researchers alike. Firstly, we begin in Section 5.1 by looking at the problem of Pareto front visualisation, which is a core component in many modern decision making workflows. We showcase how it is possible to adapt many existing visualisation strategies from the deterministic setting into the stochastic setting by appealing to our polar parametrisation result. Secondly, in Section 5.2, we illustrate how all of these Pareto front surface statistics and visualisation ideas can be used within a standard Bayesian experimental design setting. Subsequently, in Section 5.3, we highlight how our polar parametrisation result can be used within the realms of multivariate extreme value theory. Lastly, in Section 5.4, we present a Pareto front analysis of a real-world time series data set concerning the air pollution levels in North Kensington, London.
@@ 5.1 Visualisation
Visualisation is a key element in many modern multi-objective decision making pipelines. We can easily visualise the Pareto front surface for low-dimensional problems ($M\leq 3$) by using a simple two-dimensional line plot or a three-dimensional surface plot. For higher dimensional problems ($M>3$), these simple approaches no longer work and we have to become more creative when it comes to visualising the Pareto front—see the work by Tušar and Filipič (2015) for a survey on these visualisation strategies. Notably, this problem of visualising the Pareto front surface is exacerbated when we also want to include uncertainty information about the Pareto front surface distribution as well. The majority of existing work in Pareto front visualisation has largely focussed on the deterministic setting. To the best of our knowledge, very little work (if any) has focussed on the practically important problem of visualising a random Pareto front surface. In this section, we address this gap in the literature by proposing a general visualisation approach which works even in the stochastic setting. Our strategy hinges on the main result in this section (Proposition 5.1), which gives us a way to slice a high-dimensional Pareto front surface into a collection of lower-dimensional Pareto front surfaces. Conceptually, this result tells us that we can build a picture of the overall Pareto front surface by navigating the space of low-dimensional slices. To illustrate the usefulness of our new method, we built two simple dashboard applications which can be used to visualise and navigate the space of one or two-dimensional Pareto front surface slices, respectively—see Figures 10 and 11 for a snapshot of these applications.
@@ 5.1.1 Projected Pareto front surfaces
The polar parametrisation result tells us that any $M$-dimensional Pareto front is isomorphic to the set of positive unit vectors $\mathcal{S}_{+}^{M-1}$. Therefore, in theory, any strategy that can be used to visualise the set of positive unit vectors can also be used to visualise a Pareto front surface. In this section, we propose a novel idea to visualise any Pareto front surface by appealing to the following partitioning of the space of positive unit vectors:
@@
where $I=\{i_{1},\dots,i_{P}\}\subset[M]:=\{1,\dots,M\}$ is a non-empty set of $|I|=P$ ordered indices with $i_{1}<\cdots<i_{P}$; $\mathcal{V}_{+}^{M-P}:=\{\mathbf{z}\in\mathbb{R}^{M-P}_{>0}:||\mathbf{z}||_{L^{2}}<1\}$ is the set of vectors lying within the positive orthant of the $(M-P)$-dimensional sphere; and $\mathcal{P}_{I}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{P}$ is the projection mapping with $\mathcal{P}_{I}(\mathbf{a})=(a^{(i_{1})},\dots,a^{(i_{P})})\in\mathbb{R}^{P}$ for any vector $\mathbf{a}\in\mathbb{R}^{M}$. Intuitively, the partition in (25) gives us a way to slice up the set of positive unit vectors by intersecting it with the hyperplanes $\{\mathbf{y}\in\mathbb{R}^{M}:\mathcal{P}_{[M]\setminus I}(\mathbf{y})=\mathbf{v}\}$. By construction, if we project each slice onto the remaining indices, then we obtain a lower-dimensional Pareto front surface. We summarise this result in Lemma 5.1 and prove it in Section A.7. Notably, this result gives us a way to visualise the set of positive unit vectors $\mathcal{S}_{+}^{M-1}$ through its lower-dimensional projections. In the following, we will extend this visualisation strategy to work for any arbitrary Pareto front surface $A^{*}\in\mathbb{Y}_{\boldsymbol{\eta}}^{*}$.
@@
$\mathcal{P}_{I}(\phi_{I}(\mathbf{v},\boldsymbol{\lambda}))=\sqrt{1-||\mathbf{v}||_{L^{2}}^{2}}\boldsymbol{\lambda}\quad\text{ and }\quad\mathcal{P}_{[M]\setminus I}(\phi_{I}(\mathbf{v},\boldsymbol{\lambda}))=\mathbf{v}$
@@
for any polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$. In Proposition 5.1, we show that the projection of any Pareto front surface $A^{*}\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$, that is $\mathcal{P}_{I,\mathbf{v}}[A^{*}]\in\mathbb{L}_{\mathcal{P}_{I}(\boldsymbol{\eta})}$, is a valid $P$-dimensional Pareto front surface with the reference vector $\mathcal{P}_{I}(\boldsymbol{\eta})\in\mathbb{R}^{P}$. This result follows similarly in spirit to Lemma 5.1. For completeness, we explicitly state this result below in Proposition 5.1 and prove it in Section A.8. In addition, we have also included a concrete example in Figure 9, where we demonstrate how a three-dimensional Pareto front surface can be partitioned into a collection of two-dimensional slices.
@@
5.1.2 One-dimensional projection
The polar parametrisation result gives us a way to navigate a Pareto front surface by simply navigating the space of unit vectors $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$. We made a simple dashboard application to illustrate how such a navigation can be done in practice. We give a snapshot of this application in Figure 10. Below we describe the main concept of this application.
Figure 10: An illustration of the dashboard application that we made in order to navigate the space of one-dimensional projections of a Pareto front surface (Section 5.1.2).
Weights
We can navigate the space of unit vectors by moving $M$ different sliders. Each slider is associated with a weight lying in the open unit interval: $w^{(m)}\in(0,1)$ for the objectives $m=1,\dots,M$. The positive unit vector can then be constructed by normalising the weight vector appropriately: $\boldsymbol{\lambda}=\mathbf{w}/||\mathbf{w}||_{L^{2}}\in\mathcal{S}_{+}^{M-1}$.
Normalisation

### v2

@@ 4.4 Probability of domination
The probability that a set of vectors is Pareto optimal is given by the probability of domination. This probability can also be interpreted as a generalisation of the survival function for Pareto front surface distributions. By Lemma 3.1, we can write the probability of domination in term of the projected lengths
@@
for any set of vectors $Y\subseteq\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$, where $\land$ denotes the logical and operation. In practice, this probability is an intractable quantity. Nevertheless, if we can sample a collection of random Pareto front surfaces, then we can easily estimate it by using a simple Monte Carlo average. In the left plot of Figure 8, we visualise the contours associated with the estimated marginal domination probabilities for the Pareto front surface distribution described in Figure 7.
@@ 4.5 Probability of deviation
In some cases, we might be interested in computing the probability that a set of vectors lies between two potentially random Pareto front surfaces. We refer to this quantity as the probability of deviation. By using Lemma 3.1, we can write the probability of deviation in term of the projected lengths
@@
for any random Pareto front surfaces $A^{*}(\boldsymbol{\omega}),B^{*}(\boldsymbol{\omega})\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$ and any set vectors $Y\subseteq\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}]$. As with the domination probabilities in (17), this quantity can be easily estimated using a Monte Carlo average. In Figure 8, we give an illustration of the estimated marginal deviation probabilities when $A^{*}(\boldsymbol{\omega})=Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega})$ is a random Pareto front surface and $B^{*}(\boldsymbol{\omega})$ is a candidate for the expectation.
@@ 4.6 Vorob’ev statistics
Random set theory (Molchanov, 2005) extends standard concepts from Probability and Statistics to work over the space of closed sets. Some key ideas from this field have been used in earlier work in order to define some Pareto front surface statistics (Grunert da Fonseca and Fonseca, 2010; Binois et al., 2015). These works have focussed on defining probabilistic concepts over the space of domination regions $\mathcal{D}_{\preceq}:=\{\mathbb{D}_{\preceq}[A]\subset\mathbb{R}^{M}:A\subset\mathbb{R}^{M}\}$. In the following, we will review the core ideas within these works and show how they relate to our framework. To increase the generality of our discussion, we will consider working in the more general space of truncated domination regions $\mathcal{D}_{\preceq,\boldsymbol{\eta}}:=\{\mathbb{D}_{\preceq,\boldsymbol{\eta}}[A]\subset\mathbb{R}^{M}:A\subset\mathbb{R}^{M}\}$.
@@
In random set theory, one often relies on the concept of a coverage function, which gives us the probability that an element lies within a random closed set. When working over the space of truncated domination regions $\mathcal{D}_{\preceq,\boldsymbol{\eta}}$, this coverage function is simply given by the probability of domination (17): $\mathbb{P}[\mathbf{y}\in\mathbb{D}_{\preceq,\boldsymbol{\eta}}[Y^{*}_{\boldsymbol{\eta},f}]]$ for any vector $\mathbf{y}\in\mathbb{R}^{M}$. Equipped with this coverage function, we can now define many different types of probabilistic concepts by leveraging ideas from random set theory. Notably, Grunert da Fonseca and Fonseca (2010) and Binois et al. (2015) have primarily focussed on extending the Vorob’ev concepts of random set theory (Molchanov, 2005, Section 2.2), which we describe below.
@@
In addition to proposing the Vorob’ev $\alpha$-quantile, Grunert da Fonseca and Fonseca (2010) proposed using the median quantile ($\alpha=0.5$) as a potential candidate for the mean Pareto front surface. In contrast Binois et al. (2015) proposed using the Vorob’ev expectation as a candidate for the mean. The Vorob’ev expectation is defined as the $\alpha^{*}$-level Vorob’ev quantile whose hypervolume is the closest to the expected hypervolume of the random Pareto front surface. More precisely,
$\mathbb{E}^{\text{Vorob'ev}}_{\boldsymbol{\omega}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega})]:=\mathcal{Q}_{\boldsymbol{\omega}}^{\text{Vorob'ev}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega}),\alpha^{*}],$
where $\alpha^{*}$ satisfies the inequality
$\nu[\mathcal{Q}^{\text{Vorob'ev}}_{\boldsymbol{\omega}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega}),\alpha]]\leq V_{\boldsymbol{\eta},f}\leq\nu[\mathcal{Q}^{\text{Vorob'ev}}_{\boldsymbol{\omega}}[Y^{*}_{\boldsymbol{\eta},f}(\boldsymbol{\omega}),\alpha^{*}]],$ (21)
@@
Analogous to how the traditional scalar-valued expectation minimises the variance, the Vorob’ev mean is known to minimise a quantity known as the Vorob’ev deviation. Binois et al. (2015) defined the Vorob’ev deviation of a Pareto front surface distribution as the expected hypervolume of the symmetric difference between the Pareto front surface and the considered set. Equivalently, the quantity can be written as the expected hypervolume distance (9) between these two sets, that is
@@
for any measurable set $A\in\mathcal{M}(\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}])$ lying in the truncated space. As shown by Molchanov (2005, Section 2.2), the Vorob’ev mean is a minimiser to the Vorob’ev deviation over the space of measurable sets $A\in\mathcal{M}(\mathbb{D}_{\succ\mathrel{\mkern-5.0mu}\succ}[\{\boldsymbol{\eta}\}])$ subject to the condition that $\nu[A]=V_{\boldsymbol{\eta},f}$. That is,
@@
As a direct consequence of Proposition 4.2, the positive isolines of the Vorob’ev based Pareto front surfaces are all valid Pareto front surfaces because they are just instances of the quantile front (15). This result was not proven and just taken for granted in the original works, but we have shown that it follows quite naturally from our polar parameterisation and Proposition 4.2.
@@
The Vorob’ev statistics above can be estimated using the empirical quantiles as we described in Section 4.3. For the Vorob’ev mean front, we have to additionally compute the parameter $\alpha^{*}$ which satisfies the hypervolume condition (21). To approximately solve this problem, Binois et al. (2015) proposed using an iterative bisection scheme. Notably, this procedure can be very expensive because it requires evaluating the hypervolume of the quantiles several times. On the other hand, our definition of the expectation is much cheaper and simpler to estimate. In Figure 8, we present a visual comparison between the estimated Vorob’ev mean front and our sample mean front for a simple two-dimensional example. In this example, there does not appear to be any meaningful difference between these two Pareto front surfaces.
@@
4.7 A decision-theoretic construction
Many standard statistics of real-valued random variables can be constructed under the decision-theoretic perspective in which we optimise a certain scoring rule (Gneiting, 2011). Our polar parameterisation of the Pareto front surface gives us an easy way to generalise this decision-theoretic construction to the space of Pareto front surface distributions. In particular, we can recover many of the Pareto front surface statistics introduced in the previous sections as being equal to a functional $F_{\boldsymbol{\omega}}$, which minimises some expected frontier loss (8),
@@ 5 Applications
The concepts and results which we describe in this work are general and can be applied in any scenario where we want to infer or quantify properties about a random Pareto front surface. In this section, we identify a few concrete applications of these ideas which might be of interest for practitioners. Firstly, we begin in Section 5.1 by looking at the problem of Pareto front visualisation, which is a core component in many modern decision making workflows. We showcase how it is possible to adapt many existing visualisation strategies from the deterministic setting into the stochastic setting by appealing to our polar parameterisation result. Secondly, in Section 5.2, we illustrate how all of these Pareto front surface statistics and visualisation ideas can be used within a standard Bayesian experimental design setting. Subsequently, in Section 5.3, we apply our framework to extend multivariate extreme value theory to the Pareto partially ordered setting. Lastly, in Section 5.4, we present a Pareto front analysis of a real-world time series data set concerning the air pollution levels in North Kensington, London.
@@ 5.1 Visualisation
Visualisation is a key element in many modern multi-objective decision making pipelines. We can easily visualise the Pareto front surface for low-dimensional problems ($M\leq 3$) by using a simple two-dimensional line plot or a three-dimensional surface plot. For higher dimensional problems ($M>3$), these simple approaches no longer work and we have to become more creative when it comes to visualising the Pareto front—see the work by Tušar and Filipič (2015) for a survey on these visualisation strategies. Notably, this problem of visualising the Pareto front surface is exacerbated when we also want to include uncertainty information about the Pareto front surface distribution as well. The majority of existing work in Pareto front visualisation has largely focussed on the deterministic setting. To the best of our knowledge, no work has focussed on the practically important problem of visualising a random Pareto front surface. In this section, we address this gap in the literature by proposing a general visualisation approach which works even in the stochastic setting. We develop a projection mapping strategy which gives us a way to slice a high-dimensional Pareto front surface into a collection of lower-dimensional Pareto front surfaces. Conceptually, we build a picture of the overall Pareto front surface by navigating the space of low-dimensional slices. To illustrate the usefulness of our new method, we built two simple dashboard applications which can be used to visualise and navigate the space of one or two-dimensional Pareto front surface slices, respectively—see Figures 10 and 11 for a snapshot of these applications.
@@ 5.1.1 Projected Pareto front surfaces
The polar parameterisation result tells us that any $M$-dimensional Pareto front is isomorphic to the set of positive unit vectors $\mathcal{S}_{+}^{M-1}$. Therefore, in theory, any strategy that can be used to visualise the set of positive unit vectors can also be used to visualise a Pareto front surface. In this section, we propose a novel idea to visualise any Pareto front surface by appealing to the following partitioning of the space of positive unit vectors:
@@
where $I=\{i_{1},\dots,i_{P}\}\subset[M]:=\{1,\dots,M\}$ is a non-empty set of $|I|=P$ ordered indices with $i_{1}<\cdots<i_{P}$; $\mathcal{P}_{I}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{P}$ is the projection mapping with $\mathcal{P}_{I}(\mathbf{a})=(a^{(i_{1})},\dots,a^{(i_{P})})\in\mathbb{R}^{P}$ for any vector $\mathbf{a}\in\mathbb{R}^{M}$; and $\mathcal{V}_{+}^{M-P}:=\{\mathbf{z}\in\mathbb{R}^{M-P}_{>0}:||\mathbf{z}||_{L^{2}}<1\}$ is the set of vectors lying within the positive orthant of the $(M-P)$-dimensional sphere. Intuitively, the partition in (25) gives us a way to slice up the set of positive unit vectors by intersecting it with the hyperplanes $\{\mathbf{y}\in\mathbb{R}^{M}:\mathcal{P}_{[M]\setminus I}(\mathbf{y})=\mathbf{v}\}$ for $\mathbf{v}\in\mathcal{V}_{+}^{M-P}$. By construction, if we project each slice onto the remaining indices, then we obtain a lower-dimensional Pareto front surface. We summarise this result in Lemma 5.1 and prove it in Section A.7.
@@
Notably, this result gives us a way to visualise the set of positive unit vectors $\mathcal{S}_{+}^{M-1}$ through its lower-dimensional projections. In the following, we will extend this visualisation strategy to work for any arbitrary Pareto front surface $A^{*}\in\mathbb{Y}_{\boldsymbol{\eta}}^{*}$.
@@
$\mathcal{P}_{I}(\phi_{I}(\mathbf{v},\boldsymbol{\lambda}))=\sqrt{1-||\mathbf{v}||_{L^{2}}^{2}}\boldsymbol{\lambda}\quad\text{ and }\quad\mathcal{P}_{[M]\setminus I}(\phi_{I}(\mathbf{v},\boldsymbol{\lambda}))=\mathbf{v},$
@@
for any polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$. In Proposition 5.1, immediately below, we show that the projection of any Pareto front surface $A^{*}\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$, that is $\mathcal{P}_{I,\mathbf{v}}[A^{*}]\in\mathbb{L}_{\mathcal{P}_{I}(\boldsymbol{\eta})}$, is a valid $P$-dimensional Pareto front surface with the reference vector $\mathcal{P}_{I}(\boldsymbol{\eta})\in\mathbb{R}^{P}$.
@@
This result follows similarly in spirit to Lemma 5.1 and the proof is presented in Section A.8. In addition, we have also included a concrete example in Figure 9, where we demonstrate how a three-dimensional Pareto front surface can be partitioned into a collection of two-dimensional slices.
@@
5.1.2 Dashboard applications
Proposition 5.1 gives us a way to visualise an $M$-dimensional Pareto front surface distribution by navigating its lower-dimensional projections. To illustrate the utility and simplicity of this new concept, we have built two interactive dashboard applications which can be used to navigate the space of one or two dimensional slices, respectively. We present a snapshot of these applications in Figure 10 and Figure 11, respectively, for a four-dimensional Pareto front surface distribution based on the reformulated bulk carrier design problem (Tanabe and Ishibuchi, 2020). Below we elaborate on the construction of these applications in some more detail.
One-dimensional projection
To navigate the space of one-dimensional slices, we propose introducing $M$ different sliders that can be used to adjust the positive unit vectors $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. Conceptually, each slider is associated with a weight lying in the open unit interval: $w^{(m)}\in(0,1)$ for the objectives $m=1,\dots,M$. The positive unit vector can then be constructed by normalising the weight vector appropriately: $\boldsymbol{\lambda}=\mathbf{w}/||\mathbf{w}||_{L^{2}}\in\mathcal{S}_{+}^{M-1}$.